# Discovery of energy landscapes towards optimized quantum transport: Environmental effects and long-range tunneling

### v1

@@ Abstract
Carrier transport in quantum networks is governed by a variety of factors, including network dimensionality and connectivity, on-site energies, couplings between sites and whether they are short- or long-range, and environmental effects. In this work, we identify classes of quasi one-dimensional chains with energy profiles that optimize carrier transport under such influences. Specifically, we optimize on-site energies using Optax’s optimistic gradient descent and AdaMax algorithms, enabled by the JAX automatic differentiation framework. Focusing on steady-state transport, we study the system’s behavior under combined unitary and nonunitary (dephasing and dissipative) effects using the Lindblad quantum master equation. After validating our optimization scheme on short chains, we extend the study to larger systems where we identify systematic patterns in energy profiles. Our analysis reveals that different types of energy landscape enhance transport, depending on whether inter-site tunneling couplings in the chain are short- or long-range, the existence of environmental interactions, and the temperature of the environment. Our classification and insights of optimal energy landscapes offer guidance for designing efficient transport systems for electronic, photovoltaic and quantum communication applications.
@@ I Introduction
Quantum transport is ubiquitous in any non-trivial quantum system, and in the age of quantum technology must be harnessed to effectively design new devices. For example, understanding exciton transfer across networks is fundamental for light harvesting technologies, encompassing both organic solar cells [1, 2, 3, 4, 5, 6] and natural photosynthetic complexes [7, 8, 9, 10, 11]. Similarly, quantum transport of charge carriers through arrays of quantum dots [12, 13, 6, 14, 15, 16, 17, 18, 19, 20, 21] or nanoscale devices [22, 23, 24, 25, 26, 27, 28, 29, 30] plays a critical role in the development of quantum electronic devices. Also, the transmission of quantum information, whether of particles or light, across quantum networks is an increasingly active area of research, driven by the growing interest in quantum information processing. Understanding these systems is important for quantum simulations and quantum search algorithms; see, e.g., Ref. 31.
In the examples listed above, the transfer of particles and excitations is typically not purely quantum coherent: Interactions with the surrounding environment, such as phonons in the material [3], electromagnetic fields [32], or other environmental degrees of freedom [33], introduce dephasing and dissipation effects. The resultant transfer behavior is thus defined by the non-trivial confluence of coherent and incoherent effects. In practice, the transport of particles under environmental influences is often modeled using quantum walks [34, 35, 36, 37] or with a master equation [38, 39, 40, 32, 41, 42, 43, 44]. In this context, “quantum walks” [45], the analogue of classical random walks, describe the propagation of quantum particles on networks.
Optimizing transport processes in open quantum systems (OQS) remains an open and actively-investigated question. A large body of research has shown that in certain cases, for example within disordered systems, interactions with the environment leading to dynamical noise can enhance quantum transport and define an optimal region of operation. This phenomenon is known as Environmental Noise-Assisted Quantum Transport (ENAQT). While first investigated theoretically [46, 47, 48, 49, 50, 51, 52, 53], this effect was recently demonstrated in chains of trapped ions [54] and in the solid-state, within nanocrystal superlattices [55]. In fact, related turnover trends of reaction rates and electron flux with environmental interactions in the form of friction coefficients or decoherence rates are well known, and these effects have been investigated in both classical and quantum systems [56, 57, 58].
@@
A substantial body of research is focused on analyzing particular system designs, examining how a specific structure may lead to localization [70, 71, 72, 38, 73, 74], or how particular systems respond to different environmental conditions [46, 47, 48, 49, 33, 50, 51, 10, 64, 75, 76]. In contrast, relatively little effort [40, 77] has been devoted to the discovery of system configurations or energy landscapes that promote efficient transport without presupposing structural motifs.
In this paper, our goal is to discover structures that exhibit enhanced quantum transport under the combined influence of coherent and dissipative or dephasing environments, and to develop intuition into the advantage of these structures. Specifically, we focus on quasi one-dimensional tight-binding chains and consider the transfer of a single carrier from an entry site to a designated collection site. Environmental effects in the form of dephasing or thermally-induced transitions are incorporated by adding different noise terms using the Lindblad quantum master equation formalism. Figure 1 illustrates our setup: a chain of $N$ sites, with possibly long-range tunneling coupling elements between sites, leakage rate at the collection site, and environmental effects. We search for the sites’ energy profile that maximizes the outgoing flux.
@@ II Model and Dynamics
We focus on quantum transport in quasi-one-dimensional systems that are coupled to local environments (see Fig. 1). In order to focus on the essence of the problem, we make it more tractable by introducing several simplifying assumptions. First, we consider the presence of only a single carrier within the quantum system, physically corresponding to a single charge or excitation. Second, we include the environment in a phenomenological manner, using Lindblad dissipators [78]. The advantage of this description is that it applies to different environments, such as from intra- or inter-molecular vibrations [76, 79], phonons [21, 80], photons [81] or electrons [82]. We refer to our models as “quasi-one-dimensional” since we test both short-range and long-range tunneling-coupling scenarios, where in the latter case there is essentially an all-to-all connectivity in the system. In what follows, we work with units of $\hbar\equiv 1$ and $k_{B}\equiv 1$.
@@
Here, $\varepsilon_{n}$ is the energy of site n, and $J_{|n-m|}$ is the tunneling energy between site m to site n. For simplicity, these tunnelings are chosen to be real and positive. We use a power-law function to describe the tunneling energies,
@@
with $J_{max}$ as the magnitude of the nearest-neighbors tunneling. In simulations, we use $\alpha=1$ to represent long-range coupling, while $\alpha=3$ supports shorter-range tunnelings. We often refer here to the two tunneling models as the long-range tunneling model and the short-range tunneling model, referring to the magnitude of tunneling elements beyond nearest neighbors.
The carriers interact with environments— typically phonons [50, 51], photons [39, 83], local impurities [84], and other carriers [82]. Microscopically, an environment can be captured in the harmonic approximation by a collection of harmonic modes of frequency $\omega_{k}$,
@@
Here, $g_{nk}$, assumed real valued without loss of generality, are the coupling energies between the $k^{\text{th}}$ bath mode and a local operator of the system on site $n$, $\hat{L}_{n}$.
@@ II.1 OQS Model I
Here, the Lindblad jump operators are given by $|\hat{L}_{n}=\ket{n}\bra{n}$, $\Gamma_{n}$ are the dephasing rate constants, and the dynamics evolve in the site-local basis. In this case, the microscopic information from Eqs. (3) and (4) is not explicitly used, but one can associate $\Gamma_{n}$ with a microscopic model of the bath [38]. It should be noted that in the energy basis, the local dephasing model corresponds to an infinite temperature bath [85].
We assume that all sites suffer the same dephasing rate, $\Gamma_{n}\equiv\Gamma$. Regarding the range of values for $\Gamma$: theory, computations, and experiments [86, 54, 55] point out that environmental noise-assisted quantum transport is most pronounced when $\Gamma/J_{max}\sim 1$. As such, we work in this regime and set these parameters to be comparable.
@@ II.2 OQS Model II
In OQS Model II, the environment still acts locally and independently on each site, but we consider a finite-temperature bath and solve the problem in the energy basis of the system. That is, we employ the global Lindblad QME (also referred to as the exciton-basis QME in the context of exciton transport). In the global basis, transition rates between different eigenstates of the system depend on the temperature of the environment and the energy difference between those states. We write down the global Lindblad QME as
@@
where $\hat{L}_{n}=|n\rangle\langle n|$ are the site-local operators and $S_{ab}$ corresponds to the spectral function of the bath, calculated at the specific transition,
$S_{ab}=\Gamma_{0}\left|\omega_{ab}\right|\left(n_{BE}\left(|\omega_{ab}|\right)+\Theta\left(\omega_{ba}\right)\right).$ (8)
@@
Here, we assume that the baths are characterized by ohmic spectral functions. The above factor depends on the temperature of the bath; $n_{BE}(\omega)$ is the Bose-Einstein occupation factor $n_{BE}(\omega)=\frac{1}{e^{\omega/T}-1}$. Other parameters are $\Gamma_{0}$ as a dimensionless constant that dictates the strength of system-bath coupling, $\omega_{ab}=\omega_{a}-\omega_{b}$, and $\Theta(\omega)$ as the Heaviside step function. Eq. (8) imposes the detailed balance relation, thus ensuring that transitions to lower energy states are more probable than transitions to higher-energy state; if $\omega_{ab}>0$, then $S_{ab}=\Gamma_{0}\omega_{ab}\frac{e^{\omega_{ab}/T}}{e^{\omega_{ab}/T}-1}$, $S_{ba}=\Gamma_{0}\omega_{ab}\frac{1}{e^{\omega_{ab}/T}-1}$.
The dynamics imposed by Eq. (6) decouples the population and coherences by using the so-called secular approximation. It relies on the assumption that energy levels are sufficiently separated from each other such that internal coherent dynamics have timescales shorter than the overall excitation decay time.
II.3 Steady state and measure for flux
@@
In the literature, different measures have been suggested to quantify the ability of a network or a conductor to efficiently transfer carriers from an initial site to a target point. This includes the mean first passage time and other transient measures [46, 47, 54, 91, 92], as well as steady-state (or quasi-steady-state) measures, such as a rate constant and flux [50, 51, 38, 93, 94, 95, 39]. Here, we set the system under steady-state conditions, and our measure for transfer is the population flux out of a site of interest.
@@
The steady state is constructed by enforcing a constant flux of carriers entering site $1$ and leaving site $N$. These sites were chosen with the aim of achieving transport over the the longest distance allowed given a specified number of sites. Different combinations of injection and leakage sites may lead to different transport mechanisms being observed, see Refs. [54, 95]. We consider both Eqs. (5) and (6), and formally write them as $\dot{\vec{\rho}}=M\vec{\rho}$, with the density matrix written in vectorized form.
@@
To enforce the steady state, first, we add the following jump operator to the Lindblad QME (written here in the site basis), $|\hat{L}_{l}=\ket{1}\bra{N}$, along with a temperature-independent rate constant, denoted by $\gamma_{l}$. This term corresponds to a leak and injection (or trapping) process. Second, in the steady state, we have $\dot{\rho}=0$; we find the solution to this equation, $\rho_{SS}$, by replacing one of the dependent rows in $M$ (in our case, the first row) with the population normalization condition and solving $\tilde{M}\vec{\rho}^{SS}=\vec{u}$, where $\vec{u}$ is a vector of zeros besides the normalization value in the first row. $\tilde{M}$ is the modified Liouvillian matrix after adding the leakage jump operator and the normalization equation. This allows us to obtain the steady state as a matrix inversion task, rather than by obtaining the eigenvectors of $M$, which presents difficulties with automatic differentiation (see Appendix A).
@@
Our measure of transfer is the steady-state flux from the exit site $N$ of the system. It is given by
@@
The central question addressed in this work is straightforward: which families of energy profiles, $\vec{\varepsilon}=(\varepsilon_{1},\varepsilon_{2},...,\varepsilon_{N})$ maximize the flux? In particular, we investigate the role of long-range tunneling compared to short-range tunneling, and examine how environmental coupling and finite-temperature effects influence the optimal energy landscape. Given that the dynamics incorporate both coherent and incoherent processes, we anticipate that the optimal profiles will reflect the interplay between these competing mechanisms.

### v2

@@ Abstract
Carrier transport in quantum networks is governed by a variety of factors, including network dimensionality and connectivity, on-site energies, couplings between sites and whether they are short- or long-range, leakage processes, and environmental effects. In this work, we identify classes of quasi-one-dimensional chains with energy profiles that optimize carrier transport under such influences. Specifically, we optimize on-site energies using Optax’s optimistic gradient descent and AdaMax algorithms, enabled by the JAX automatic differentiation framework. Focusing on nonequilibrium steady-state transport, we study the system’s behavior under combined unitary and nonunitary (dephasing and dissipative) effects using the Lindblad quantum master equation. After validating our optimization scheme on short chains, we extend the study to larger systems where we identify systematic patterns in energy profiles. Our analysis reveals that different types of energy landscapes enhance transport, depending on whether inter-site tunneling couplings in the chain are short- or long-range, the existence of environmental interactions, and the temperature of the environment. Our classification and insights of optimal energy landscapes offer guidance for designing efficient transport systems for electronic, photovoltaic and quantum communication applications.
@@ I Introduction
Quantum transport is ubiquitous in any non-trivial quantum system, and in the age of quantum technology must be harnessed to effectively design new devices. For example, understanding exciton transfer across networks is fundamental for light harvesting technologies, encompassing both organic solar cells [21, 40, 101, 90, 14, 48] and natural photosynthetic complexes [31, 87, 88, 103, 120]. Similarly, quantum transport of charge carriers through arrays of quantum dots [66, 114, 48, 10, 5, 20, 108, 2, 72, 15, 34] or nanoscale devices [93, 35, 7, 119, 109, 81, 6, 4, 78] plays a critical role in the development of quantum electronic devices. Also, the transmission of quantum information, whether of particles or light, across quantum networks is an increasingly active area of research, driven by the growing interest in quantum information processing. Understanding these systems is important for quantum simulations and quantum search algorithms; see, e.g., Ref. 99.
In the examples listed above, the transfer of particles and excitations is typically not purely quantum coherent: interactions with the surrounding environment, such as phonons in the material [101, 67], electromagnetic fields [25], or other environmental degrees of freedom, introduce dephasing and dissipation effects. The resultant transfer behavior is thus defined by the non-trivial confluence of coherent and incoherent effects. In practice, the transport of particles under environmental influences is often modeled using quantum walks [82, 42, 68, 58] or with a master equation [18, 19, 24, 25, 94, 95, 47, 59]. In this context, “quantum walks” [107], the analogue of classical random walks, describe the propagation of quantum particles on networks.
Optimizing transport processes in open quantum systems (OQS) remains an open and actively-investigated question. In this work, our goal is to use modern optimization techniques to discover energetic landscapes that promote transport without imposing a priori structural assumptions. Our analysis focuses on one-dimensional chains, exploring both short-range and long-range tunneling regimes, and considering the effects of dephasing and thermal relaxation.
A large body of research has shown that in certain cases, for example within disordered systems, interactions with the environment leading to dynamical noise can enhance quantum transport and define an optimal region of operation. This phenomenon is known as Environmental Noise-Assisted Quantum Transport (ENAQT). While first investigated theoretically [75, 79, 12, 16, 17, 111, 73, 74, 104, 115, 116, 22, 60], this effect was recently demonstrated in chains of trapped ions [70], and in the solid-state within nanocrystal superlattices [8]. In fact, related turnover trends of reaction rates and electron flux with environmental interactions in the form of friction coefficients or decoherence rates are well known, and these effects have been investigated in both classical and quantum systems [46, 26, 100].
@@
A substantial body of research is focused on analyzing particular system designs, examining how a specific structure may lead to localization [77, 39, 84, 18, 28, 51], or how particular systems respond to different environmental conditions, including optimization of the surrounding environments towards maximal efficiency [75, 11, 85, 74, 104, 67, 115, 116, 73, 103, 49, 63, 91]. In contrast, relatively little effort [69, 23] has been devoted to the discovery of system configurations or energy landscapes that promote efficient transport without presupposing structural motifs, despite the identification of the energetic landscape as the primary factor for transport efficiency [24].
In this paper, our goal is to discover structures that exhibit enhanced quantum transport under the combined influence of coherent and dephasing or dissipative environments, and to develop intuition into the advantage of these structures. Specifically, we focus on quasi one-dimensional tight-binding chains in which inter-site tunneling couplings are short- or long-range and optimize the chains’ energy landscape towards transfer of carriers from an entry site to a designated collection site. Environmental effects in the form of dephasing or thermally-induced transitions are incorporated by adding different noise terms using the Lindblad quantum master equation formalism. While prior studies employed a forward approach, fixing the system’s energy structure and then analyzing the resulting behavior, as was done in studies of photosynthetic proteins [75, 79, 12, 16], we adopt an inverse method, searching for system energy configurations that optimize transport efficiency.
Figure 1 illustrates our setup: a chain of $N$ sites, with possibly long-range tunneling coupling elements between sites, leakage rate at the collection site, and environmental effects. We search for the sites’ energy profile that maximizes the outgoing flux.
@@ II Model and Dynamics
We focus on quantum transport in quasi-one-dimensional systems that are coupled to local environments (see Fig. 1). In order to focus on the essence of the problem, we make it more tractable by introducing several simplifying assumptions. First, we consider the presence of only a single carrier within the quantum system, physically corresponding to a single charge or excitation. Second, we include the environment in a phenomenological way, using Lindblad dissipators [9]. The advantage of this description is that it applies to different environments, such as intra- or intermolecular vibrations [91, 53], phonons [34, 50], photons [43] or electrons [89]. We refer to our models as “quasi-one-dimensional” since we test both short-range and long-range tunneling coupling scenarios, where in the latter case there is essentially an all-to-all connectivity in the system. In what follows, we work with units of $\hbar\equiv 1$ and $k_{B}\equiv 1$.
@@
Here, $\varepsilon_{n}$ is the energy of site n, and $J_{|n-m|}$ is the tunneling energy between site m to site n. For simplicity, these tunneling elements are chosen to be real and positive. We use a power-law function to describe the tunneling energies,
@@
with $J_{max}$ as the magnitude of the nearest-neighbors tunneling coupling. In simulations, we use $\alpha=1$ to represent long-range coupling, while $\alpha=3$ supports shorter-range tunnelings. Throughout this work, we distinguish between two tunneling models: the long-range model, characterized by significant tunneling beyond nearest neighbors, and the short-range model, where tunneling is effectively limited to nearest neighbors.
The carriers interact with environments — typically phonons, local impurities, and other carriers. Microscopically, an environment can be captured in the harmonic approximation by a collection of harmonic modes of frequency $\omega_{k}$,
@@
Here, $g_{nk}$, taken to be real-valued without loss of generality, denotes the coupling energy between the $k^{\text{th}}$ bath mode and a local system operator $\hat{L}_{n}$ acting on site $n$.
@@ II.1 OQS Model I
Here, the Lindblad jump operators are given by $|\hat{L}_{n}=\ket{n}\bra{n}$, $\Gamma_{n}$ are the dephasing rate constants, and the dynamics evolve in the site-local basis. In this case, the microscopic information from Eqs. (3) and (4) is not explicitly used, but one can associate $\Gamma_{n}$ with a microscopic model of the bath [18]. It should be noted that in the energy basis, the local dephasing model corresponds to an unstructured infinite temperature bath. This is because, for asymptotically small $\gamma_{l}$, the steady state solution of the dynamics corresponds to the completely mixed state with equal populations at each site, independent of energy profile [18, 105]. A nonzero value for $\gamma_{l}$ affects the population distribution due to the nonequilibrium setting, thus showing deviations from the maximally mixed state. However, it can be readily proved that the nonequilibrium distribution is nonthermal, even if $\Gamma_{n}$ is made site- and temperature-dependent.
We assume that all sites suffer a comparable dephasing and set $\Gamma_{n}\equiv\Gamma$. Theory, simulations, and experiments [38, 70, 8] indicate that noise-assisted quantum transport is most pronounced when $\Gamma/J_{max}\sim 1$. Accordingly, we focus on this regime and choose parameters of comparable magnitude.
@@ II.2 OQS Model II
In OQS Model II, the environment still acts locally and independently on each site, but we consider a finite-temperature bath and solve the problem in the energy basis of the system; that is, we employ the global Lindblad QME (also referred to as the exciton-basis QME in the context of exciton transport). In the global basis, transition rates between different eigenstates of the system depend on the temperature of the environment and the energy difference between those states. We write the global Lindblad QME as
@@
where $\hat{L}_{n}=|n\rangle\langle n|$ are the site-local operators and $S_{ab}$ corresponds to the bath spectral function [25], calculated at the specific transition,
$S_{ab}=\Gamma_{0}\left|\omega_{ab}\right|\left(n_{BE}\left(|\omega_{ab}|\right)+\Theta\left(\omega_{ab}\right)\right).$ (8)
@@
Here, we assume that the baths are characterized by ohmic spectral functions, but one can readily select other forms. The above factor depends on the temperature of the bath; $n_{BE}(\omega)$ is the Bose-Einstein occupation factor $n_{BE}(\omega)=\frac{1}{e^{\omega/T}-1}$. Other parameters are $\Gamma_{0}$ as a dimensionless constant that dictates the strength of system-bath coupling, $\omega_{ab}=\omega_{a}-\omega_{b}$, and $\Theta(\omega)$ as the Heaviside step function. Eq. (8) imposes the detailed balance relation, thus ensuring that transitions to lower energy states are more probable than transitions to higher-energy state; if $\omega_{ab}>0$, then $S_{ab}=\Gamma_{0}\omega_{ab}\frac{e^{\omega_{ab}/T}}{e^{\omega_{ab}/T}-1}$, $S_{ba}=\Gamma_{0}\omega_{ab}\frac{1}{e^{\omega_{ab}/T}-1}$. The dynamics imposed by Eq. (6) decouples the population and coherences by using the so-called secular approximation. It relies on the assumption that energy levels are sufficiently separated from each other such that internal coherent dynamics have timescales shorter than the overall excitation decay time.
II.3 Nonequilibrium steady state and measures for flux
@@
In the literature, different measures have been suggested to quantify the ability of a network or a conductor to efficiently transfer carriers from an initial site to a target point. This includes the mean first passage time and other transient measures [75, 85, 70, 106, 117], as well as steady-state (or quasi-steady-state) measures, such as a rate constant and flux [115, 116, 18, 54, 32, 33, 19]. Here, we set the system under nonequilibrium steady-state conditions, and our measure for transfer is the population flux out of site $N$.
@@
The nonequilibrium steady state is constructed by enforcing a constant flux of carriers entering site $1$ and leaving site $N$. These sites were chosen with the aim of achieving transport over the longest distance allowed given a specified number of sites. Different combinations of injection and leakage sites may lead to different transport mechanisms being observed, see Refs. [70, 33]. We consider both Eqs. (5) and (6), and formally write them as $\dot{\vec{\rho}}=M\vec{\rho}$, with the density matrix written in vectorized form following the methods described in Ref. [62].
@@
To enforce the nonequilibrium steady state, first, we add the following jump operator to the Lindblad QME (written here in the site basis), $|\hat{L}_{l}=\ket{1}\bra{N}$, along with a temperature-independent rate constant, denoted by $\gamma_{l}$. This term corresponds to a leak and injection (or trapping) process. Second, in the steady state, we have $\dot{\rho}=0$; we find the solution to this equation, $\rho_{SS}$, by replacing one of the dependent rows in $M$ (in our case, the first row) with the population normalization condition and solving $\tilde{M}\vec{\rho}^{SS}=\vec{u}$, where $\vec{u}$ is a vector of zeros besides the normalization value in the first row. $\tilde{M}$ is the modified Liouvillian matrix after adding the leakage jump operator and the normalization equation. This allows us to obtain the steady state as a matrix inversion task, rather than by obtaining the eigenvectors of $M$, which presents difficulties with automatic differentiation (see Appendix A).
@@
Our measure of transfer is the nonequilibrium steady-state flux from the exit site $N$ of the system. It is given by
@@
The flux measures the rate of population leakage from the exit site $N$. Since the jump operators that we use to describe the environment do not deplete population, the flux is constant along the chain, though for nonlocal tunneling, it has to be calculated along all contributing bonds [67].
We now make several comments on the nonequilibrium steady state transport framework, the measure we use, and its relation to other measures.
First, we know that when $\gamma_{l}=0$, and for a nonzero $\Gamma$, the system reaches in the long time limit its equilibrium steady state solution, which for OQS Model I is the completely mixed state. For $\gamma_{l}\neq 0$, the fixed point is a nonequilibrium steady state.
Second, several studies have employed transient measures to quantify transfer efficiency. In particular, Ref. [70] used the following definition, $\tau_{T}=\frac{1}{T}\int_{0}^{T}p_{N}(\tau)d\tau$, with $p_{N}(t)$ as the time-evolving population of the target site, $N$. The integration time $T$ must be chosen carefully: it has to be chosen such that time evolution is long enough to build population in the target site from the initial condition, and short enough to be distinct from the final equilibrium state. For a chain with population initialized at site 1 with the target site $N$, a common choice is $T=N/J_{max}$, where $J_{max}$ is the nearest neighbor tunneling energy and $N$ the site-to-site distance between initial to target sites [70]. However, this choice of integration time can significantly affect results. Moreover, the expression for $T$ must generally be re-evaluated and adjusted for different energy profiles and dephasing strengths. Overall, time-dependent measures of this kind may lack robustness, and varying $T$ influences the inferred transport efficiency.
Third, we note that different measures correspond to different experimental settings. Experiments may be performed as an initial-condition problem, where an excitation is prepared at an entry site and the transfer efficiency is quantified by the rate or probability of arrival at a designated target site. In contrast, the nonequilibrium steady-state framework corresponds to a boundary-condition problem, analogous to a current–voltage setup, where a constant drive or input maintains a steady current across the system. These two scenarios, imposing initial conditions versus enforcing nonequilibrium boundary conditions, can generally lead to different trends. This does not imply that either approach is incorrect; rather, it highlights that they describe distinct experimental settings. In Ref. [54], we compared transient and nonequilibrium steady state measures and demonstrated cases where the two approaches yield consistent results, as well as situations in which they diverge.